OpenStudy (umulas):

Quick trig question: For a question, I have : Find the exact value cos(-7pi/12); what do I do? Should I do this cos{pi/6-3pi/4] cosxcosy+sinxsiny √3/2 * -√2/2 = √3-√2/4 + 1/2*√2/2 √3 --- 4 Can anyone explain what I did wrong?

OpenStudy (umulas):

The answers I have is √6 + 2 √6/4 - √2/4 √2 - √6 √2/4 - √6/4

OpenStudy (jdoe0001):

well, 1/6 - 3/4 = -1/12; \(\bf \ne \) -7/12

OpenStudy (umulas):

Hmmm? 1/6 = 2/12 3/4 = 9/12 2 - 9 -7 ---- = -- 12 12

OpenStudy (jdoe0001):

I can see -4/12 - 3/12 = -7/12

OpenStudy (jdoe0001):

ohhh yea... hmmm, you're right, I spoke too soon, I forgot a small piece on the numerator

hartnn (hartnn):

cosx cosy+sinx siny = (√3/2) * (-√2/2) + (1/2) (√2/2) = -√6/4 + √2/4

OpenStudy (umulas):

Ahhhhhh; I forgot to multiply the square roots :P.

hartnn (hartnn):

yeah .. :)